# Measuring cosmic baryon density with FRB and GW data

### v1

@@ 2.2 GW standard siren data
We use 47 GW standard sirens with SNRs greater than 11 and an inverse false alarm rate exceeding 4 yr from GWTC-3, covering a redshift range of approximately $z<0.8$. The sample consists of 42 BBHs, three neutron star-black holes (NSBHs), and two binary neutron stars (BNSs).
We obtain the redshift of GW170817 from the electromagnetic (EM) counterpart (Abbott et al., 2017a, b, c). For GW events without EM counterparts, we obtain their redshifts through the dark siren method (Schutz, 1986; Mastrogiovanni et al., 2023; Gray et al., 2023) and incorporate the GLADE+ catalog (Dálya et al., 2022). In the following, we provide a brief introduction to the likelihood function of GW standard sirens. We denote the source parameters of each GW event by $\bm{\theta}$, and the population-level parameters by $\bm{\Lambda}$, including population model parameters and cosmological parameters. For GW events without EM counterparts, assuming a “scale-free” prior on the expected number of detected events $N_{\exp}$, i.e., $\pi(N_{\exp})\propto 1/N_{\exp}$, the likelihood function of the GW data set $\{x_{\rm GW}\}$ composed of $N_{\rm GW}$ GW events (Mastrogiovanni et al., 2021; Abbott et al., 2023b), is given by
$\mathcal{L}_{\rm GW}(\{x_{\rm GW}\}|\bm{\Lambda})\propto\prod_{i}^{N_{\rm GW}}\frac{\int{\rm d}\bm{\theta}\mathcal{L}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})\frac{{\rm d}N_{\rm CBC}(\bm{\Lambda})}{{\rm d}\bm{\theta}{\rm d}t}}{\int{\rm d}\bm{\theta}P_{\rm det}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})\frac{{\rm d}N_{\rm CBC}(\bm{\Lambda})}{{\rm d}\bm{\theta}{\rm d}t}},$ (8)
where $\mathcal{L}_{\rm GW}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})$ is the parameter estimation likelihood, quantifying the uncertainties in measuring the GW source parameters $\bm{\theta}$ through waveform analysis. ${\rm d}N_{\rm CBC}(\bm{\Lambda})/{\rm d}\bm{\theta}{\rm d}t$ is the CBC detection rate in the detector frame, which includes population models of GW sources, cosmological models, and redshift information from the GLADE+ galaxy catalog. In assessing the completeness of the GLADE+ catalog, we adopt a Schechter function to describe the $B$-band absolute magnitude distribution of galaxies (Kochanek et al., 2001). We further assume that the probability of a galaxy being the host of the GW source is proportional to its $B$-band luminosity. $P_{\rm det}(x_{{\rm GW},i}|\bm{\theta},\bm{\Lambda})$ is the detection probability of a GW event with parameters of $\bm{\theta}$ and $\bm{\Lambda}$. The denominator of Eq. (8) accounts for the GW selection effect.
@@
For GW dark sirens, population models can help infer redshift for the sources, especially in scenarios where the galaxy catalog is significantly incomplete (Taylor et al., 2012; Ezquiaga & Holz, 2022; Mastrogiovanni et al., 2023; Gray et al., 2023). We adopt the same population model assumptions as Abbott et al. (2023a), but restrict our analysis to the Power Law + Peak BBH mass model, which is best supported by the data. Considering additional models would increase computational cost and lies beyond the scope of this work. For BNSs, we assume that neutron star masses are uniformly distributed in the range of 1–3 $M_{\odot}$. In the case of NSBH systems, the black hole component is drawn from the same distribution as the primary black holes in BBHs, while the neutron star component follows the mass distribution assumed for BNSs. The redshift distribution of GW sources is modeled using a phenomenological prescription that traces the cosmic star formation rate. Consistent with Abbott et al. (2023a), we disregard potential redshift evolution in the BBH mass distribution as well as variations in the GW detection rate arising from the spin distribution. These effects are currently subdominant compared to the statistical uncertainties present in the existing GW observations.
In cases where EM counterparts are detected, the information from EM observations should be incorporated alongside GW data. Since GW and EM detections are statistically independent, the overall likelihood is given by
$\mathcal{L}_{\rm EM+GW}(x_{{\rm GW},i}|\bm{\theta})\propto\mathcal{L}_{\rm EM}(x_{{\rm GW},i}|z,\Omega)\mathcal{L}_{\rm GW}(x_{{\rm GW},i}|\bm{\theta}),$ (9)
where $\mathcal{L}_{\rm EM}(x_{{\rm GW},i}|z,\Omega)$ quantifies the measurement uncertainties of the redshift $z$ and the sky localization $\Omega$ from EM observations. We assume that the probability of having EM observations is 1 once an EM counterpart is present (Gray et al., 2020; Mastrogiovanni et al., 2024). This assumption holds only within the current GW detection range, as it remains much smaller than that of EM observations.
2.3 Joint analysis of FRB and GW data
@@
Figure 1: Normalized posterior distributions of $\Omega_{\rm b}~{}$from 104 FRBs (orange) and 104 FRBs + 47 GWs (indigo). The median and $1\sigma$ credible interval from FRB + GW is $\Omega_{\rm b}=0.0488\pm 0.0064$, marginalized over $f_{\rm d}$ and empirical DM model parameters. The errorbars in different colors and linestyles are shown under different assumptions: FRB + $H_{0}$ (Planck18 prior) and FRB + $H_{0}$ (SH0ES prior) adopt Gaussian priors on $H_{0}$ from Planck18 (Aghanim et al., 2020) and SH0ES (Riess et al., 2022), respectively; FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW adopt priors on empirical DM model parameters from the IllustrisTNG simulation (Zhang et al., 2020, 2021) and on $f_{\rm d}$ from the redshift-dependent form of Macquart20 (Macquart et al., 2020), respectively. The blue, gray, and red shaded bands show the results from CMB + BBN (Cooke et al., 2018), Macquart20, and Yang22 (Yang et al., 2022), respectively. Note that Yang22 adopts the TNG prior, a flat prior on $H_{0}$in $\mathcal{U}(67,73)$ km s-1 Mpc-1, and $f_{\rm d}=0.82$. All results are at the $1\sigma$ level.
@@
We jointly constrain cosmological and FRB-related DM parameters, and adopt flat priors for the following seven parameters: $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ (Yang et al., 2022), $\Omega_{\rm b}\in\mathcal{U}(0.01,0.10)$, $H_{0}\in\mathcal{U}(20,140)~{}\rm km\ s^{-1}\ Mpc^{-1}$, $f_{\rm d}\in\mathcal{U}(0.84,0.96)$, [^2] This prior on $f_{\rm d}$ is informed by observations of the baryon fraction in stars ($f_{*}$) and cold gas ($f_{\rm cold}$), independent of FRBs. Assuming a conservative flat prior for $f_{*}$ extended by a factor of five, the diffuse baryon fraction is inferred via $f_{\rm d}=1-f_{*}-f_{\rm cold}$ (Connor et al., 2025). $F\in\mathcal{U}(0.011,1.0)$, $\mu_{\rm host}\in\mathcal{U}(3.0,5.3)$, and $\sigma_{\rm host}\in\mathcal{U}(0.2,2.0)$ (Macquart et al., 2020). The GW population model parameters include $\alpha$, $\beta$, $m_{\rm min}$, $m_{\rm max}$, $\Delta m$, $\mu_{\rm g}$, $\sigma_{\rm g}$, $\lambda_{\rm peak}$, $\gamma$, $\kappa$, and $z_{\rm p}$, which are fixed to the inferred values of Abbott et al. (2023a).
@@
We utilize the bilby [^3] https://github.com/bilby-dev/bilby package to perform Markov Chain Monto Carlo (MCMC) analysis and choose the emcee [^4] https://emcee.readthedocs.io/en/stable sampler. We use the ICAROGW [^5] https://github.com/simone-mastrogiovanni/icarogw package to do the hierarchic Bayesian analysis of GW standard sirens, and the GetDist [^6] https://github.com/cmbant/getdist package to plot the marginalized posterior distributions. The final likelihood $\mathcal{L}$ used in our analysis is the combination of $\mathcal{L}_{\mathrm{FRB}}$ and $\mathcal{L}_{\rm GW}$ in Eqs. (6) and (8), respectively, which is written as
@@
Table 1: $1\sigma$ constraints on cosmological parameters $\Omega_{\rm b}$ and $H_{0}$.
@@
3 Results and discussions
Figure 2: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm b}$ and $\Omega_{\rm b}h_{70}$ versus $H_{0}$, $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$, using FRB (orange) and FRB + GW (purple) data. The correlation coefficient $r$ is labeled in each panel to quantify parameter degeneracies.
@@ 3 Results and discussions
In this section, we report and discuss the constraints on $\Omega_{\rm b}~{}$in Table 1. The posteriors for $\Omega_{\rm b}~{}$and the corresponding credible intervals are shown in Figure 1. Using FRB data alone yields a relatively weak constraint of $\Omega_{\mathrm{b}}=0.052_{-0.028}^{+0.012}$ at the $1\sigma$ level. Incorporating GW standard sirens significantly narrows this to $\Omega_{\mathrm{b}}=0.0488\pm 0.0064$, by breaking the strong degeneracy between $\Omega_{\mathrm{b}}$ and $H_{0}$ via the independent $H_{0}$ measurement provided by GWs. Remarkably, this late-Universe estimate exhibits concordance with early-Universe values from BBN and CMB at the $\sim 13\%$ level, linking the gap between the early and late Universe. Note that a similarly precise constraint is simultaneously obtained for the Hubble constant, $H_{0}=71.6^{+4.4}_{-8.6}~{}\mathrm{kms^{-1}~{}Mpc^{-1}}$, primarily driven by the GW dataset alone, which gives $H_{0}=72.6^{+5.2}_{-8.9}~{}\mathrm{kms^{-1}~{}Mpc^{-1}}$. More data is required to arbitrate the $H_{0}$ tension. A detailed discussion on $H_{0}$ and DM parameters is provided in Appendix A.
@@
We compare our $\Omega_{\rm b}~{}$results with those from previous works. Macquart20 and Yang22 constrained $\Omega_{\rm b}~{}$using different samples of localized FRBs under external assumptions of $H_{0}=70$ km s-1  Mpc-1and a flat prior on $H_{0}\in\mathcal{U}(67,73)$ km s-1  Mpc-1, as shown by the gray and red bands in Figure 1, respectively. The constraint from Yang22 compares favorably to ours, since the current GW standard siren sample remains too limited to outperform their adopted $H_{0}$ prior. Applying the external $H_{0}$ priors from Planck18 or SH0ES to our FRB sample alone yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively; the corresponding error bars are shown in Figure 1. This suggests that the $H_{0}$ tension would lead to two biased $\Omega_{\rm b}~{}$estimates. Although the induced “$\Omega_{\rm b}$ inconsistency” in our sample is within $1\sigma$ uncertainty ($\sim 0.95\sigma$), it may become non-negligible with larger samples in updated DM analyses. Nevertheless, the joint FRB + GW result assumes no prior on $H_{0}$ and instead provides a late-Universe measurement of baryons, independent of both CMB and local distance ladder. This FRB + GW synergy may establish a novel paradigm for the cosmic baryon census, analogous to CMB + BBN, and serve as a powerful low-redshift cosmological probe in the era of next-generation FRB and GW surveys (Zhang et al., 2023).
@@
We also investigate potential improvements to our results. FRB cosmology is subject to uncontrolled systematics mainly from modeling $\mathrm{DM}_{\mathrm{IGM}}$ and $\mathrm{DM}_{\mathrm{host}}$ distributions, as well as the weakly constrained $f_{\rm d}$, which affects the mean value of $\mathrm{DM}_{\mathrm{IGM}}$. To reduce these systematics, cosmological simulations are commonly used to calibrate these DM parameters ($F$, $\mu_{\rm host}$, $\sigma_{\rm host}$, and $f_{\rm d}$) at different redshifts. Accordingly, we consider two cases with simulation-based priors: (i) fixing the empirical DM model parameters to the best-fit results from the IllustrisTNG simulation (hereafter TNG) (Zhang et al., 2020, 2021); and (ii) further fixing $f_{\rm d}$ to a redshift-dependent form following Macquart20, as implemented in the public FRB code [^7] https://github.com/FRBs/FRB. The constraints from FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW are $\Omega_{\mathrm{b}}=0.0491\pm 0.0054$ and $\Omega_{\mathrm{b}}=0.0507\pm 0.0052$, respectively, with the corresponding errorbars shown in Figure 1. Both results are consistent with the FRB + GW constraint within the $1\sigma$ uncertainty, suggesting that the simulation-based priors can improve current constraints without introducing significant bias. However, such agreement at the current level does not ensure the absence of systematic errors when applied to larger future FRB and GW samples, in which systematic errors may become dominant. Therefore, a detailed assessment of how the DM parameters affect the inferred baryon density is of great importance.
@@
To assess the impact of systematic uncertainties on $\Omega_{\mathrm{b}}$ and verify the robustness of our constraint, we analyze the parameter correlations using the FRB + GW dataset. The lower row of Figure 2 shows the marginalized posteriors of $\Omega_{\rm b}~{}$and other relevant parameters, with correlation coefficients $r$ labeled in each panel to quantify parameter degeneracies.
@@
On the one hand, the degeneracy between $\Omega_{\rm b}$ and $H_{0}$ is the strongest among all parameter pairs ($r=-0.73$), implying that the inferred value of $\Omega_{\rm b}$ is highly sensitive to how $H_{0}$ is determined. While FRB + $H_{0}$ analyses rely on external priors, which are controversial due to the $H_{0}$ tension, FRB + GW analyses rely on GW data to provide $H_{0}$, which is uniquely derived from self-calibrating absolute distance measurements. Therefore, the $\Omega_{\rm b}$ constraint from FRB + GW is robust against the $H_{0}$-related bias and free from assumptions affected by the $H_{0}$ tension.
@@
On the other hand, the correlations between $\Omega_{\rm b}~{}$and DM parameters are moderate. These DM parameters are not properly constrained, both in this work (see Appendix A for details) and earlier studies (James et al., 2022), and thus act as nuisance parameters that bias the inferred $\Omega_{\rm b}$. In the joint FRB + GW analysis, the correlation coefficients between $\Omega_{\rm b}$ and $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$ are $-0.28$, $0.40$, $-0.40$, and $0.17$, respectively. For comparison, the upper row of Figure 2 shows the constraint on $\Omega_{\rm b}h_{70}$, where $h_{70}=H_{0}/(70~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}})$. This parameter absorbs the strong degeneracy, enabling it to be effectively constrained using FRB data alone (Macquart et al., 2020; Connor et al., 2025). In this case, the corresponding correlations are notably stronger: $-0.41$, $0.57$, $-0.62$, and $0.27$, respectively. This comparison suggests that directly measuring $\Omega_{\rm b}$ reduces degeneracies with the nuisance parameters and is more robust against the associated systematic uncertainties than $\Omega_{\rm b}h_{70}$.
@@
We conclude that incorporating GW data into the analysis of FRB dispersion measures enables a direct and effective measurement of $\Omega_{\rm b}$ and, more importantly, avoids biases from the $H_{0}$ tension and mitigates those associated with nuisance model parameters. This demonstrates that the synergy between FRBs and GWs provides a robust estimate of the cosmic baryon density.
4 Conclusion

### v2

@@ II.2 GW standard siren data
We use 47 GW standard sirens with SNRs greater than 11 and an inverse false alarm rate exceeding 4 yr from GWTC-3222https://gwosc.org/GWTC-3, covering a redshift range of approximately $z<0.8$. The sample consists of 42 BBHs, three neutron star-black holes (NSBHs), and two BNSs. We do not consider GW events detected in the LVK O4a run, mainly because during this run, Virgo is not online, and all GW events are detected by two LIGOs, resulting in relatively poorer sky localizations and luminosity distance measurements LIGOScientific:2025jau.
We obtain the redshift of GW170817 from the electromagnetic (EM) counterpart LIGOScientific:2017adf; LIGOScientific:2017vwq; LIGOScientific:2017zic. For GW events without EM counterparts, we infer their redshifts through the dark siren method Schutz:1986gp; Mastrogiovanni:2023emh; Gray:2023wgj and incorporate the GLADE+ catalog Dalya:2021ewn. In the dark siren method, assuming the GW sources all originate from galaxies, we can cross-match sky localizations of GW events with galaxies collected in galaxy catalogs, and identify a serious potential host galaxy. These potential host galaxies provide redshift priors for GW events, which help to infer cosmological parameters in a hierarchical Bayesian framework. Furthermore, GW population models can help break degeneracies of masses and redshifts in GW detections, thereby providing additional redshift information Taylor:2011fs; Ezquiaga:2022zkx; Mastrogiovanni:2023emh; Gray:2023wgj. In this paper, we use the public pipeline ICAROGW Mastrogiovanni:2023zbw to compute the GW likelihood function, denoted as $\mathcal{L}_{\rm GW}$. The explicit formulas can be found in Refs Mastrogiovanni:2023emh; Gray:2023wgj.
@@
We adopt the same population model assumptions as Ref. LIGOScientific:2021aug, but restrict our analysis to the Power Law + Peak BBH mass model, which is best supported by the GWTC-3 data. Considering additional models would increase computational cost and lies beyond the scope of this work. For BNSs, we assume that neutron star masses are uniformly distributed in the range of 1–3 $M_{\odot}$. In the case of NSBH systems, the black hole component is drawn from the same distribution as the primary black holes in BBHs, while the neutron star component follows the mass distribution assumed for BNSs. The redshift distribution of GW sources is modeled using a phenomenological prescription that traces the cosmic star formation rate. Consistent with Ref. LIGOScientific:2021aug, we disregard potential redshift evolution in the BBH mass distribution as well as variations in the GW detection rate arising from the spin distribution. These effects are currently subdominant compared to the statistical uncertainties present in the existing GW observations.
II.3 Joint analysis of FRB and GW data
@@
Figure 1: Normalized posterior distributions of $\Omega_{\rm b}$~from 104 FRBs (orange) and 104 FRBs + 47 GWs (indigo). The median and $1\sigma$ credible interval from FRB + GW is $\Omega_{\rm b}=0.0488\pm 0.0064$, marginalized over $f_{\rm d}$ and empirical DM model parameters. The errorbars in different colors and linestyles are shown under different assumptions: FRB + $H_{0}$ (Planck18 prior) and FRB + $H_{0}$ (SH0ES prior) adopt Gaussian priors on $H_{0}$ from Planck18 2020Planck and SH0ES Riess:2021jrx, respectively; FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW adopt priors on empirical DM model parameters from the IllustrisTNG simulation Zhang:2020mgq; Zhang:2020xoc and on $f_{\rm d}$ from the redshift-dependent form of Macquart20 Macquart:2020lln, respectively. The blue, gray, and red shaded bands show the results from CMB + BBN Cooke:2017cwo, Macquart20, and Yang22 Yang:2022ftm, respectively. Note that Yang22 adopts the TNG prior, a flat prior on $H_{0}$ $\in$ $\mathcal{U}(67,73)$ km s-1 Mpc-1, and $f_{\rm d}=0.82$. All results are at the $1\sigma$ level.
@@
We jointly constrain cosmological and FRB-related DM parameters, and adopt flat priors for the following seven parameters: $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ Yang:2022ftm, $\Omega_{\rm b}\in\mathcal{U}(0.01,0.10)$, $H_{0}\in\mathcal{U}(20,140)~\rm km\ s^{-1}\ Mpc^{-1}$, $f_{\rm d}\in\mathcal{U}(0.84,0.96)$, [^3] This prior on $f_{\rm d}$ is informed by observations of the baryon fraction in stars ($f_{*}$) and cold gas ($f_{\rm cold}$), independent of FRBs. Assuming a conservative flat prior for $f_{*}$ extended by a factor of five, the diffuse baryon fraction is inferred via $f_{\rm d}=1-f_{*}-f_{\rm cold}$ connor2025gas. $F\in\mathcal{U}(0.011,1.0)$, $\mu_{\rm host}\in\mathcal{U}(3.0,5.3)$, and $\sigma_{\rm host}\in\mathcal{U}(0.2,2.0)$ Macquart:2020lln. For the GW population model parameters, we consider two scenarios. In the first, following Ref. LIGOScientific:2021aug, we fix all population-model parameters to their fiducial values. In the second, we allow the BBH mass-distribution parameters and the redshift-distribution parameters of GW sources to vary freely, adopting the same prior ranges as the spectral siren analysis in Ref. LIGOScientific:2021aug, to assess the impact of population-model uncertainties on our results. This treatment is similar to Ref. Gray:2023wgj; however, unlike Ref. LIGOScientific:2025jau, we do not free the mass-distribution parameters of neutron-star sources.
@@
We utilize the bilby [^4] https://github.com/bilby-dev/bilby package to perform Markov Chain Monte Carlo (MCMC) analysis and choose the emcee [^5] https://emcee.readthedocs.io/en/stable sampler. The GetDist [^6] https://github.com/cmbant/getdist package to plot the marginalized posterior distributions. The final likelihood $\mathcal{L}$ used in our analysis is the combination of $\mathcal{L}_{\mathrm{FRB}}$ and $\mathcal{L}_{\rm GW}$, which is written as
@@
Table 1: $1\sigma$ constraints on cosmological parameters $\Omega_{\rm b}$ and $H_{0}$. The results are derived from $104$ localized FRBs alone or in combination with $47$ GW events, under different assumptions. For all cases, the $1\sigma$ constraints on DM parameters are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$. Unless otherwise specified, the FRB + GW case assumes a flat prior of $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ and fixed GW population parameters. Note that $H_{0}$ is in units of km s-1  Mpc-1.
@@
    FRB + GW (GW parameters free)     $0.0448^{+0.0094}_{-0.0076}$     $79.9^{+5.0}_{-17.0}$
III Results and discussions
In this section, we first report the constraints on the baryon density as our main results in Sec. III.1, and then extend the analysis to assess the impacts of FRB and GW population modeling assumptions in Secs. III.2 and III.3, respectively.
III.1 Main results
@@
First, we report the main results for the baryon-density estimate in the first block of Table 1. The posteriors for $\Omega_{\rm b}$~and the corresponding credible intervals are shown in Fig. 1. Using FRB data alone yields a relatively weak constraint of $\Omega_{\mathrm{b}}=0.052_{-0.028}^{+0.012}$ at the $1\sigma$ level. Incorporating GW standard sirens significantly narrows this to $\Omega_{\mathrm{b}}=0.0488\pm 0.0064$, by breaking the strong degeneracy between $\Omega_{\mathrm{b}}$ and $H_{0}$ via the independent $H_{0}$ measurement provided by GWs. Remarkably, this late-Universe estimate exhibits concordance with early-Universe values from BBN and CMB at the $\sim 13\%$ level, linking the gap between the early and late Universe. Note that a similarly precise constraint is simultaneously obtained for the Hubble constant, $H_{0}=71.6^{+4.4}_{-8.6}~\mathrm{kms^{-1}~Mpc^{-1}}$, primarily driven by the GW dataset alone, which gives $H_{0}=72.6^{+5.2}_{-8.9}~\mathrm{kms^{-1}~Mpc^{-1}}$. More data is required to arbitrate the $H_{0}$ tension. A detailed discussion on $H_{0}$ and DM parameters is provided in Appendix A.
@@
We compare our $\Omega_{\rm b}$~results with those from previous works. Macquart20 and Yang22 constrained $\Omega_{\rm b}$~using different samples of localized FRBs under external assumptions of $H_{0}=70$ km s-1  Mpc-1and a flat prior on $H_{0}\in\mathcal{U}(67,73)$ km s-1  Mpc-1, as shown by the gray and red bands in Fig. 1, respectively. The constraint from Yang22 compares favorably to ours, since the current GW standard siren sample remains too limited to outperform their adopted $H_{0}$ prior. Applying the external $H_{0}$ priors from Planck18 or SH0ES to our FRB sample alone yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively; the corresponding error bars are shown in Fig. 1. This suggests that the $H_{0}$ tension would lead to two biased $\Omega_{\rm b}$~estimates. Although the induced $\Omega_{\rm b}$ inconsistency” in our sample is within $1\sigma$ uncertainty ($\sim 0.95\sigma$), it may become non-negligible with larger samples in updated DM analyses. Nevertheless, the joint FRB + GW result assumes no prior on $H_{0}$ and instead provides a late-Universe measurement of baryons, independent of both CMB and local distance ladder. This FRB + GW synergy may establish a novel paradigm for the cosmic baryon census, analogous to CMB + BBN, and serve as a powerful low-redshift cosmological probe in the era of next-generation FRB and GW surveys Zhang:2023gye.
We also assess the dependence on the prior for $\Omega_{\rm m}$. Relaxing the prior from $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ to a broad flat prior $\Omega_{\rm m}\in\mathcal{U}(0,1)$ yields $\Omega_{\rm b}=0.0454^{+0.0066}_{-0.0082}$ and $H_{0}=72.7^{+4.7}_{-9.0}$ km s-1  Mpc-1, as shown in the third row of Table 1. Compared with the fiducial case ($\Omega_{\rm b}=0.0488\pm 0.0064$), the slightly lower value of $\Omega_{\rm b}$ is driven by the biased constraint on $\Omega_{\rm m}$ under the broad prior, $\Omega_{\rm m}=0.225^{+0.060}_{-0.180}$, which propagates through the positive correlation between $\Omega_{\rm b}$~and $\Omega_{\rm m}$. Therefore, we caveat that our present analysis moderately depends on the $\Omega_{\rm m}$ prior, which can be improved with larger future FRB and GW samples to better constrain $\Omega_{\rm m}$.
III.2 Impact of FRB model parameters
Figure 2: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm b}$ and $\Omega_{\rm b}h_{70}$ versus $H_{0}$, $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$, using FRB (orange) and FRB + GW (indigo) data. The correlation coefficient $r$ is labeled in each panel to quantify parameter degeneracies. Note that $H_{0}$ is in units of km s-1  Mpc-1.
@@ III.2 Impact of FRB model parameters
Next, we assess the impact of FRB population modeling on the baryon census, as listed in the second block of Table 1. In our fiducial analysis in Sec. III.1, no specific FRB population model is imposed and all DM parameters are left free to avoid potential bias. To explore how such modeling affects the $\Omega_{\rm b}$ inference, we apply several commonly used FRB population models and evaluate their potential to improve the precision of $\Omega_{\rm b}$. FRB cosmology is subject to uncontrolled systematics mainly from modeling $\mathrm{DM}_{\mathrm{IGM}}$ and $\mathrm{DM}_{\mathrm{host}}$ distributions, as well as the weakly constrained $f_{\rm d}$, which affects the mean value of $\mathrm{DM}_{\mathrm{IGM}}$. To reduce these systematics, cosmological simulations are commonly used to calibrate these DM parameters ($F$, $\mu_{\rm host}$, $\sigma_{\rm host}$, and $f_{\rm d}$) at different redshifts. Accordingly, we consider two cases with simulation-based priors: (i) fixing the empirical DM model parameters to the best-fit results from the IllustrisTNG simulation (hereafter TNG) Zhang:2020mgq; Zhang:2020xoc; and (ii) further fixing $f_{\rm d}$ to a redshift-dependent form following Macquart20, as implemented in the public FRB code [^7] https://github.com/FRBs/FRB. The constraints from FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW are $\Omega_{\mathrm{b}}=0.0491\pm 0.0054$ and $\Omega_{\mathrm{b}}=0.0507\pm 0.0052$, respectively, with the corresponding errorbars shown in Fig. 1. Both results are consistent with the FRB + GW constraint within the $1\sigma$ uncertainty, suggesting that the simulation-based priors can improve current constraints without introducing significant bias. However, such agreement at the current level does not ensure the absence of systematic errors when applied to larger future FRB and GW samples, in which systematic errors may become dominant. Therefore, a detailed assessment of how the DM parameters affect the inferred baryon density is of great importance.
@@
To assess the impact of systematic uncertainties on $\Omega_{\mathrm{b}}$ and verify the robustness of our constraint, we analyze the parameter correlations using the FRB + GW dataset. The lower row of Fig. 2 shows the marginalized posteriors of $\Omega_{\rm b}$~and other relevant parameters, with correlation coefficients $r$ labeled in each panel to quantify parameter degeneracies.
@@
On the one hand, the degeneracy between $\Omega_{\rm b}$ and $H_{0}$ is the strongest among all parameter pairs (r=0.73r=-0.73), implying that the inferred value of $\Omega_{\rm b}$ is highly sensitive to how $H_{0}$ is determined. While FRB + $H_{0}$ analyses rely on external priors, which are controversial due to the $H_{0}$ tension, FRB + GW analyses rely on GW data to provide $H_{0}$, which is uniquely derived from self-calibrating absolute distance measurements. Therefore, the $\Omega_{\rm b}$ constraint from FRB + GW is robust against the $H_{0}$-related bias and free from assumptions affected by the $H_{0}$ tension.
@@
On the other hand, the correlations between $\Omega_{\rm b}$~and DM parameters are moderate. These DM parameters are not properly constrained, both in this work (see Appendix A for details) and earlier studies James:2022dcx, and thus act as nuisance parameters that bias the inferred $\Omega_{\rm b}$. In the joint FRB + GW analysis, the correlation coefficients between $\Omega_{\rm b}$ and $f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$ are 0.28-0.28, $0.40$, 0.40-0.40, and $0.17$, respectively. For comparison, the upper row of Fig. 2 shows the constraint on $\Omega_{\rm b}h_{70}$, where $h_{70}=H_{0}/(70~\mathrm{km~s^{-1}~Mpc^{-1}})$. This parameter absorbs the strong degeneracy, enabling it to be effectively constrained using FRB data alone Macquart:2020lln; connor2025gas. In this case, the corresponding correlations are notably stronger: 0.41-0.41, $0.57$, 0.62-0.62, and $0.27$, respectively. This comparison suggests that directly measuring $\Omega_{\rm b}$ reduces degeneracies with the nuisance parameters and is more robust against the associated systematic uncertainties than $\Omega_{\rm b}h_{70}$.
@@
We conclude that the impact of FRB population modeling on the inferred $\Omega_{\rm b}$ is negligible within our analysis. The results also show that, within our FRB + GW framework, incorporating GW data into the analysis of FRB dispersion measures enables a direct and effective measurement of $\Omega_{\rm b}$ and, more importantly, avoids biases from the $H_{0}$ tension and mitigates those associated with nuisance model parameters. This demonstrates that the synergy between FRBs and GWs provides a robust estimate of the cosmic baryon density.
III.3 Impact of GW model parameters
Figure 3: Posterior distributions ($1\sigma$ and $2\sigma$ credible regions) for $\Omega_{\rm m}$, $H_{0}$, and the typical GW population parameter $\mu_{g}$ using FRB + GW data. Two cases are shown: GW parameters fixed within the PowerLaw + Peak model (indigo) and the parameters allowed to vary freely and marginalised over (gray). The dotted lines indicate the median values of each posterior, except for $\mu_{g}$ in the GW-parameter-fixed case, where the dotted line marks the prior value $\mu_{g}=32.27$ adopted following Ref. LIGOScientific:2021aug. Note that $H_{0}$ is in units of km s-1  Mpc-1.
Last, we assess the impact of GW population modeling on our baryon census, as listed in the third block of Table 1. In our fiducial analysis in Sec. III.1, no FRB population model is imposed, whereas the PowerLaw + Peak GW population model is adopted for the GW population, as is commonly used in GW-based cosmological inference LIGOScientific:2021aug. To investigate the robustness of our $\Omega_{\rm b}$~estimate against this assumption, we release the GW population model parameters and marginalize over them. We find that joint FRB + GW analysis with GW model parameters free yields a higher value of $H_{0}=79.9^{+5.0}_{-17.0}$ km s-1  Mpc-1and a correspondingly lower value of $\Omega_{\rm b}=0.0448^{+0.0094}_{-0.0076}$. The $\Omega_{\rm b}$ result ($\sim 20\%$ precision) remains consistent with the CMB + BBN determination at the $1\sigma$ level.
To illustrate how the GW population prior affects $\Omega_{\rm b}$, we focus on the parameter $\mu_{g}$, which denotes the mean of the Gaussian component in the primary mass distribution within the PowerLaw + Peak model. This parameter is moderately correlated with the inferred $H_{0}$ from GW observations LIGOScientific:2021aug, making it suitable for accounting for the impact of population modeling. As shown in Fig. 3, $\mu_{g}$ exhibits a negative degeneracy with $H_{0}$ and a corresponding positive degeneracy with $\Omega_{\rm b}$. Consequently, compared with the prior value $\mu_{g}=32.27$ adopted in the GW–parameter-fixed (indigo) case, the lower inferred value of $\mu_{g}=31.5$ in the GW–parameter-free (gray) case leads to a higher $H_{0}$ and thus a lower $\Omega_{\rm b}$. Thus, within our FRB + GW framework, the inferred $\Omega_{\rm b}$ is more sensitive to GW-related parameters than to FRB-related ones (see Sec. III.2). It is because $H_{0}$, the essential parameter to infer $\Omega_{\rm b}$ in our framework, is more strongly degenerate with the GW population parameters than with the FRB-related parameters.
We conclude that, to obtain an unbiased and robust estimate of $\Omega_{\rm b}$, caution is required when adopting GW-population assumptions, due to their strong impact on the inferred $H_{0}$. Nevertheless, our test without imposing any GW population model reinforces that our fiducial result obtained with the PowerLaw + Peak model is reliable and not significantly biased.
IV Conclusion